OpenStudy (anonymous):

Prove n!/k!(n-k+1)!+n!/k!(n-k)!=(n+1)!/k!(n+1-k)!

OpenStudy (anonymous):

\[\frac{ n! }{ k!(n-k+1)! }+\frac{ n! }{ k!(n-k)! }=\frac{ (n+1)! }{ k!(n+1-k)! }\]

OpenStudy (abb0t):

n!/k!*(n-k)! + n!/(k-1)!*(n-k+1)! =n!/k(k-1)!(n-k)!+n!/(k-1)!(n-k+1)(n-k… =[n!/(k-1)!(n-k)!]{1/k+1/(n-k+1)} =[n!/(k-1)!(n-k)!][(n+1)/k(n-k+1)] =(n+1)!/k!(n-k+1)!